OpenStudy (anonymous):

Could you help me with this next question? If 4m-3/7=3, which is the value? a.) 6 b.) 10.5 c.) 37 d.) 68.5

OpenStudy (anonymous):

@piperkors

OpenStudy (anonymous):

\(\Large 4m-\frac{3}{7}=3\\\Large 4m=3+\frac{3}{7}=\frac{24}{7}\\\Large m=\frac{24}{7} \div4=\frac{24}{7\times4}=\frac{6}{7}\)